Logiweb dictionary of problemone

Up Help

0 0 problemone
1 0 pred calc
2 0 pc1
3 0 pc2
4 0 pc3
5 0 pc4
6 0 pc5
7 0 pc6
8 0 pc7
9 0 pc8
10 0 pc9
11 0 pc10
12 0 pc11
13 0 pc12
14 0 pcmp
15 0 pcded
16 0 pcia
17 0 pcie
18 0 pcdeduction
19 0 trivia
20 0 iatest
21 0 andintro
22 0 andelim1
23 0 andelim2
24 0 orintro1
25 0 orintro2
26 0 orelim
27 0 notintro
28 0 notnotintro

29 0 notnotelim
30 0 mt
31 0 pbc
32 0 repeat
33 0 lem

34 2 ∗ ≡ ∗ 35 2 ∗ = ∗ 36 1 ¬∗ 37 2 ∗ ∧ ∗ 38 2 ∗ ∨ ∗ 39 2 ∀ ∗ . (∗) 40 2 ∃ ∗ . (∗) 41 2 ∗ ∈ ∗

The pyk compiler, version 0.grue.20060417+ by Klaus Grue GRD-2006-07-07.UTC:10:03:47.052758 = MJD-53923.TAI:10:04:20.052758 = LGT-4658983460052758e-6